OpenStudy (anonymous):

Use the discriminant to find the number of real solutions of the following quadratic equation: 2x2 + 2x + 2 = 0. A. no real solutions B. 1 real solution C. 2 real solutions D. 3 real solutions

OpenStudy (anonymous):

2(x^2+x+1)

OpenStudy (anonymous):

=0

OpenStudy (anonymous):

so one solution?

OpenStudy (anonymous):

2(x+1)^2=0

OpenStudy (anonymous):

Divide both sides with 2.

OpenStudy (anonymous):

(x+1)^2=0

OpenStudy (anonymous):

x+1=0

OpenStudy (anonymous):

x=-1

OpenStudy (anonymous):

i have no clue what you are trying to do. it askes for solutions not evaluate

OpenStudy (anonymous):

It IS the solution.

OpenStudy (anonymous):

THE solution.

OpenStudy (anonymous):

since its -1 that means no real solution

OpenStudy (anonymous):

It is real and negative.

OpenStudy (anonymous):

so ONE solution?

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Real_number

OpenStudy (anonymous):

I think yes. If you want to be sure goe on this webpage http://www.wolframalpha.com/

OpenStudy (anonymous):

It helps a lot when you need help for equations.